OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11).

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

what is the slope of given line?

OpenStudy (anonymous):

can u find?

OpenStudy (anonymous):

y=-4x-2?

OpenStudy (anonymous):

compare y=mx+c m is slope now what is m?

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

the slope is -4?

OpenStudy (anonymous):

right now pependicular line has slope = negative invese m of perpe =-1/m what is -1/m?

OpenStudy (anonymous):

-1/-4?

OpenStudy (anonymous):

=1/4 is slope points are given plug in equ y-y1=m(x-x1) what will be equ?

OpenStudy (anonymous):

4-(-11)=-4(-1-(-16)?

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

no y-(-11)=1/4(x-(-16)

OpenStudy (anonymous):

y+11=1/4x+4 y=1/4x-7

OpenStudy (anonymous):

so how do i write everything out lol

OpenStudy (anonymous):

you trid but dont put y and x in equ x1 and y1 are given

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

slope of given line =-4 slope of perpendicular line =-1/m=1/4 equ of perpen having slope 1/4 and passing through (-16, -11). is y-(-11)=1/4(x-(-16) y+11=1/4x+4 y=1/4x-7